OpenStudy (anonymous):

Please help me on this statics problem. I don't know what I am doing wrong here!

OpenStudy (anonymous):

wait, im posting question

OpenStudy (anonymous):

the second picture is my attempt at the problem

OpenStudy (anonymous):

i need help on problem 1. I tried using the equilibirum equations of ridged bodies...

OpenStudy (perl):

its difficult to read your problem

OpenStudy (anonymous):

@perl is this more clearer?

OpenStudy (anonymous):

Help? anyone?

OpenStudy (anonymous):

... your solution seems fine, but since this is static and equilibrium problem, summation of forces in vertical is zero, same with horizontal... summation of moments at any point should be zero also...

OpenStudy (anonymous):

... the reaction at roller D must be against the roller... horizontal only, no vertical component...

OpenStudy (anonymous):

in the 1st question, you have 3 unknowns, cable tensions \(T_B\text{ and }T_C\) and roller reaction \(R_D\)...and the 700kg mass is equivalent to a weight of 6867N...\[+\uparrow\sum F_y=0\]\[-6867+\frac{4}{5}T_B-\frac{4}{\sqrt{20}}T_C=0~~~~~(1)\]

OpenStudy (anonymous):

\[+\rightarrow\sum F_x=0\]\[\frac{3}{5}T_B+\frac{2}{\sqrt{20}}T_C-R_D=0~~~~~(2)\]

OpenStudy (anonymous):

\[+cw \sum M_D=0\]\[-6867(6)+T_B(3)-T_C(2)=0~~~~~(3)\]

OpenStudy (anonymous):

can work on equations (1) and (3) to solve tensions \(T_B\text{ and }T_C\)... then equation (2) for \(R_D\)

OpenStudy (vincent-lyon.fr):

There must be an infinity of solutions to your problem, since there are two cables which can be more or less taut without changing the equilibrium.

OpenStudy (anonymous):

@Orion1213 i took the moment about c cause wouldn't the horizontal and vertical forces of the tension cancel out at D if i took the moment there since they pass through D?

OpenStudy (micahm):

what the problem

OpenStudy (micahm):

what the name of the lesson

OpenStudy (micahm):

this is what www.physicsclassroom.com/ says When all the forces that act upon an object are balanced, then the object is said to be in ... That is, all horizontal components must add to 0 Newton and all vertical ... Suppose the tension in both of the cables is measured to be 50 N

OpenStudy (anonymous):

@micahm i posted the picture to the problem. I know that but i dont think im doing something right here. I have the equations and such but when i plug in the values, my denominator is equal to zero...

OpenStudy (micahm):

the first picture at the top

OpenStudy (micahm):

i have a site the might help you have you ever use http://www.freemathhelp.com/

OpenStudy (anonymous):

The site you gave me is for math...